# Automated Model Tuning for Multifidelity Uncertainty Propagation in Trajectory Simulation

### v1

@@ 2.1 Pilot Sampling
It can be seen in Equation (7) that evaluating the estimator variance (and thus solving ACV sample allocation and model tuning problems) requires knowledge of correlations among the models, $\underline{\rho}(\underline{\beta})$ and $\mathbf{P}(\underline{\beta})$. In general, these correlations must be estimated empirically by drawing $N_{pilot}$ random samples of $\underline{\mathbf{z}}$ and evaluating each model, i.e., by using pilot sampling.
@@
There are three different approaches, or solution modes available in Dakota[31], for solving ACV sample allocation optimization with respect to pilot sampling: 1) offline pilot, 2) pilot projection, and 3) online pilot. The offline pilot approach treats pilot sampling as an offline cost that precedes an online phase where the available budget is allocated. That is, the initial cost of evaluating the pilot samples is excluded from the budget. Note that this is the most commonly used approach in the multifidelity uncertainty quantification (UQ) literature [e.g., 15, 10, 16]. This mode can be employed to obtain an “oracle” reference solution since it provides access to a potentially fully converged set of statistics, without impacting the available computational resources within $C_{budget}$.
@@
Pilot projection, by contrast, assumes that the pilot samples will be reused and thus includes their cost within the overall computational budget. To do so, the sample allocation optimization problem is constrained to have the pilot investment as a lower bound. Typically, $N_{pilot}$ is chosen such that the pilot evaluation is a modest proportion of the overall computational budget, then projects the performance that an ACV estimator would obtain based on the initial model correlation estimates. That is, the estimator variance is minimized over $\{N,N_{i}\}$, but the correlations are not updated, the additional sample evaluations are not carried out, and therefore the final moment statistics are not computed.
@@
Online pilot, or iterated ACV [29, 31], mode begins in the same manner as pilot projection, using a (typically small) initial pilot evaluation to project estimator performance. However, this approach then aims to iterate towards the optimal sample allocation [^3] In term of the number of samples shared with the HF model. while continuously constraining the sample profile to reuse the current collection of pilot samples. It does so by repeating the following steps: 1) compute the optimal sample allocation given the current estimate of model correlations, 2) calculate the change in number of shared samples, $\Delta N$, between the current and previous sample allocation, 3) evaluate a new $\Delta N$ pilot samples [^4] In practice, $\Delta N$ may be decreased by an under-relaxation parameter to avoid executing unnecessary model evaluations due to inaccurate model statistics estimates. and update the estimate of models’ correlations. The process stops when $\Delta N\leq 0$.
@@
Each of these three solution modes excel in different settings. Offline pilot is commonly used to generate reference or oracle solutions that result from (approximately) exact a priori knowledge of model statistics. Pilot projection is helpful for evaluating the performance of different ACV estimators or for selecting the best ensemble of models. Online pilot mode has the advantage of balancing accuracy and cost in an online analysis, helping to avoid under-estimating $N_{pilot}$ (incurring reduced accuracy) or over-estimating $N_{pilot}$ (incurring reduced efficiency). Note that multiple approaches can be combined in a single analysis, e.g., using pilot projection to select the best ACV method and then online pilot to iterate on it. As shown later, all three solution modes are used for different purposes in this work.
@@ 2.2 Optimization Approach
Notably, Equation (7) illustrates that model tuning (optimizing $\underline{\beta}$) is tied with sample allocation optimization (optimizing $\underline{\mathbf{z}}$). In this work, a bi-level optimization strategy is chosen to solve Equation (1) where $\underline{\beta}$ is optimized on the outer loop with the inner loop performing sample allocation optimization for each instance of $\underline{\beta}$:
@@
Here, it is emphasized that the inner loop optimization problem depends on $\underline{\beta}$ through the required estimates of model correlations $\underline{\rho}(\underline{\beta})$ and $\mathbf{P}(\underline{\beta})$. The bi-level approach decouples the model tuning problem, allowing different off-the-shelf optimization algorithms for the outer level to be combined with varying ACV estimator schemes (Table 1) on the inner level. This includes the parametric ACV approach for an additional search over optimal model hierarchies on the inner level for added flexibility. Note that it is also possible to treat this as a single all-at-once optimization; although, this latter approach would require the construction of a surrogate for each correlation term in Eq. (7) as a function of the hyperparameters $\underline{\beta}$ (see [20, 31]).
@@
The primary challenge of the bi-level approach is that the outer-loop objective function is noisy since $\mathbb{V}ar\!\left[\tilde{Q}\right$] is evaluated from estimates of $\underline{\rho}(\underline{\beta})$ and $\mathbf{P}(\underline{\beta})$ based on finite (and often limited) pilot sampling. Note that either online pilot or pilot projection can be used to solve the inner-level sample allocation optimization depending on whether the accuracy or efficiency of the solution will be prioritized, respectively. Pilot projection is adopted in this work in order to avoid over-investment in the model tuning process, since only the samples for $\underline{\beta}^{*}$ can be reused
@@
For the outer loop, the efficient global optimization (EGO) algorithm [32] is used for its ability to handle the effects of numerical noise while minimizing $\mathbb{V}ar\!\left[\tilde{Q}\right$] . EGO is a surrogate-based optimizer that approximates the objective function with a Gaussian process (GP) response surface and then iteratively refines the GP using points selected from the maximization of an expected improvement (EI) acquisition function. A strength of EGO is its ability to select the next candidate point in such a way that balances the need to minimize the approximate objective (exploitation) with the need to improve the approximation where it is most uncertain (exploration). Algorithm 1 summarizes the use of EGO for solving the model tuning optimization problem.
@@
In this work, the portion of the overall computational budget, $C_{budget}$, dedicated to model tuning can be loosely controlled by specifying the number of pilot samples, $N_{pilot}$, and maximum number of iterations, $N_{iter}$ for the EGO algorithm. It is important to point out that a potentially significant portion of computation performed during model tuning can be reused in a subsequent multifidelity UQ analysis. For example, $N_{pilot}$ evaluations of the high-fidelity model can be saved and reused as well as the evaluations of the low-fidelity models at the optimal hyperparameters, $\underline{\beta}^{*}$. Note that using a convergence tolerance that terminates EGO based on diminishing returns from model tuning could be used in lieu of (or in addition to) setting $N_{iter}$. Doing so would have some practical advantages, but was not pursued here and instead left for future studies.
@@
where it is again assumed that all $M$ low-fidelity models are tunable. Thus, when performing a subsequent multifidelity UQ analysis, the overall computational budget is modified as $C_{budget}-C_{\mathcal{T}}$.
@@
1:Input: Initial hyperparameter dataset $\mathcal{D}_{0}=\{(\underline{\beta}_{i},J_{i})\}_{i=1}^{N_{init}}$, number of iterations $N_{iter}$, number of pilot samples $N_{pilot}$, computational budget $C_{budget}$
2:Let: $J(\underline{\beta})=\min_{\underline{\mathbf{z}}}\mathbb{V}ar\!\left[\tilde{Q}(\underline{\beta},\underline{\mathbf{z}})\right]\quad\text{s.t.}\quad\mathcal{C}(\underline{\beta},\underline{\mathbf{z}})<C_{budget}$ $\triangleright$ EGO objective function
3:for $t=1$ to $N_{iter}$ do
4:  Fit Gaussian Process surrogate model to $\mathcal{D}_{t-1}$
5:  Estimate $\underline{\rho}(\underline{\beta}_{t})$ and $\mathbf{P}(\underline{\beta}_{t})$ with $N_{pilot}$ pilot samples
6:  Select $\underline{\beta}_{t}$ to maximize the expected improvement in $J$
7:  Update dataset: $\mathcal{D}_{t}=\mathcal{D}_{t-1}\cup\{(\underline{\beta}_{t},J_{t})\}$
8:end for
[…]
10:Return: Optimal hyperparameters $\underline{\beta}^{*}$
2.3 Summary: Online ACV Estimator with Model Tuning
Figure 1 summarizes the overall procedure for obtaining ACV estimators, $\tilde{Q}\left(\underline{\beta},\underline{\alpha},\underline{\mathbf{z}}\right)$, via Equation (5) in the practical setting where neither optimal low-fidelity model settings ($\underline{\beta}$) nor knowledge of model statistics ($\mathbf{P}$, $\underline{\rho}$) are available a priori. See Algorithm 2 for more details.
Figure 1: Overview of procedure for online multifidelity UQ with model tuning to build ACV estimators without a priori knowledge of optimal low-fidelity model settings or model statistics.
@@ 2.3 Summary: Online ACV Estimator with Model Tuning
Algorithm 2 Online ACV Estimator with Automated Model Tuning
[…]
4:Bounds for unknown low-fidelity hyperparameters
[…]
@@
7:Solve model tuning optimization problem $\triangleright$ Algorithm 1 Require: $N_{iter}$, $N_{pilot}$, bounds for $\underline{\beta}$ Ensure: $\underline{\beta}^{*}$ Ensure: $C_{\mathcal{T}}$ $\triangleright$ Model tuning overhead
8:
9:Perform online ACV sample allocation optimization Require: $\underline{\beta}^{*}$ Require: $C_{budget}-C_{\mathcal{T}}$ $\triangleright$ Remaining ACV computational budget Ensure: $\underline{\mathbf{z}}^{*}$ $\triangleright$ Optimal sample allocation Ensure: $\mathbf{P}$, $\underline{\rho}$ $\triangleright$ Final estimates of model statistics Ensure: $\mathbb{V}ar\!\left[\tilde{Q}\left(\underline{\beta}^{*},\underline{\mathbf{z}}^{*}\right)\right]$ $\triangleright$ Estimated variance Note: $N_{pilot}$ pilot samples are available from previous step and need not be recomputed.
@@
11:Evaluate models according to sample allocation Require: $\underline{\mathbf{z}}^{*}$, $\underline{\beta}^{*}$ Ensure: Outputs for all models evaluated on $\underline{\mathbf{z}}^{*}$. Note: Shared pilot samples in $\underline{\mathbf{z}}^{*}$ are available from previous step and need not be recomputed.
12:
13:Form ACV estimator $\triangleright$ Equation (5) Require: Sample allocation model outputs, model statistics ($\mathbf{P}$, $\underline{\rho}$) Ensure: $\tilde{Q}$

### v2

@@ 2.4 Pilot Sampling
It can be seen from Equation (12) that solving Eq. 10 for the combined ACV sample allocation and model tuning requires knowledge of the dependence of model correlations with respect to the hyperparameters, $\underline{\rho}(\underline{\beta})$ and $\mathbf{P}(\underline{\beta})$. In general, these correlations must be estimated empirically (for fixed hyperparameter values) by drawing $N_{pilot}$ random samples of $\underline{\mathbf{z}}$ and evaluating each model, i.e., by using pilot sampling. Additional strategies to capture the dependence on $\beta$ should be introduced, as detailed in the next section.
@@
Historically, there are three different solution modes available in Dakota [28] for solving ACV sample allocation optimization with respect to pilot sampling: 1) offline pilot, 2) pilot projection, and 3) online pilot. The offline pilot approach treats pilot sampling as an offline cost that precedes an online phase where the available budget is allocated. That is, the initial cost of evaluating the pilot samples is excluded from the budget. Note that this is the most commonly used approach in the multifidelity uncertainty propagation literature [e.g., 15, 10, 16]. This mode can be employed to obtain an “oracle” reference solution since it provides access to a potentially fully converged set of correlations/covariances, without impacting the available computational resources within $C_{budget}$.
@@
Pilot projection, by contrast, is an online mode that assumes that the pilot samples will be reused and thus includes their cost within the overall computational budget. To do so, the sample allocation optimization problem is constrained to have the pilot investment as a lower bound. Typically, $N_{pilot}$ is chosen such that the pilot evaluation is a modest proportion of the overall computational budget, and then the performance of an ACV estimator is projected based on the initial model correlation estimates. That is, the estimator variance is minimized over $\{N,N_{i}\}$, but the pilot-based correlations are not updated, the additional sample evaluations are not carried out, and therefore the final moment statistics are not computed.
@@
Online pilot mode, or iterated ACV [33, 28], begins in the same manner as pilot projection, using a (typically small) initial pilot evaluation to project estimator performance. However, this approach then aims to iterate towards the optimal sample allocation in term of the number of shared samples evaluated for all models, since these sample increments can be used to update the correlation estimates. The iteration proceeds by repeating the following steps: 1) compute the optimal sample allocation given the current estimate of model correlations, 2) calculate the change in number of shared samples, $\Delta N$, between the current and previous sample allocation, 3) evaluate a new $\Delta N$ pilot samples [^5] In practice, initial $\Delta N$ estimates may be damped using an under-relaxation parameter to mitigate potential over-estimation of shared samples when model correlations are inaccurate. and update the estimated model correlations. The process stops when $\Delta N\leq 0$, then transitioning to the evaluation of low-fidelity sample increments $N_{i}$ followed by the roll-up of final statistics.
@@
Each of these three solution modes are useful in different settings. Offline pilot mode is commonly used to generate reference or oracle solutions that result from (approximately) exact a priori knowledge of model statistics. Pilot projection mode is helpful for evaluating the performance of different ACV estimators or for selecting the best ensemble of models. Online pilot mode has the advantage of balancing accuracy and cost in an online analysis, helping to avoid under-estimating $N_{pilot}$ (incurring reduced accuracy) or over-estimating $N_{pilot}$ (incurring reduced efficiency). Note that multiple approaches can be combined in a single analysis, e.g., using pilot projection to select the best ACV method and then online pilot mode to iterate with this selection. As shown later, all three solution modes are used for different purposes in this work [^6] Starting with version 6.20, Dakota adds a fourth solution mode for offline projection, augmenting the existing (online) pilot projection in order to support all combinations of online vs. offline and pilot vs. projection. This mode will not be included in the studies that follow..
@@ 2.5 Optimization Approach
The numerical solution of the optimization problem described in Equation (10) is now discussed: this equation illustrates that model tuning (optimizing $\underline{\beta}$) is tied with sample allocation optimization (optimizing $\underline{\mathbf{z}}$). Two solutions strategies can be designed for this task [28], namely an all-at-once or a bi-level strategy. These two strategies offer different trade-offs, and are described below. In the case of an all-at-once optimization, a surrogate for each term that depends on $\underline{\beta}$ is needed. In particular, since the dependency of $\mathbb{V}ar\!\left[\tilde{Q}\left(\underline{\mathbf{z}},\underline{\beta}\right)\right]$ on $\underline{\beta}$ (Equation (10)) is encapsulated in Equation12, it is necessary to obtain a surrogate for each term of $\underline{\rho}$ and $\mathbf{P}$ which, accounting for the symmetry in $\mathbf{P}$, corresponds to a total number of $M(M+1)/2$ surrogates. Moreover, for each of the $M$ low-fidelity models, a surrogate to represent the dependency of its cost $w_{i}$ on $\beta_{i}$ is needed. Each of these surrogates can potentially have a different dimension since each $\beta_{i}$ vector has a dimension associated with the hyperparameters of the $i$th model. To build these surrogates, the pilot sampling solution described in the previous section can be extended to a set of points in the support of $\underline{\beta}$, e.g., by using a regular tessellation. At each location $\underline{\beta}^{b}$ of the tesselation of the hyperparameter space, a standard pilot sampling procedure is carried out and all the required statistics are evaluated, i.e., $\underline{\rho}(\underline{\beta}^{b})$, $\mathbf{P}(\underline{\beta}^{b})$, and $w_{i}(\underline{\beta}^{b})$. Surrogates are then built from $N_{b}$ realizations of these quantities $\underline{\rho}(\underline{\beta}^{b})$, $\mathbf{P}(\underline{\beta}^{b})$, and $w_{i}(\underline{\beta}^{b})$ for $b=1,\dots,N_{b}$ via standard techniques like regression or interpolation. Once the surrogates are trained, Equation (10) can be solved by resorting to standard solutions for the multifidelity resources allocations. For instance, an example of this strategy based on local quadratic polynomials and Sequential Least Squares Programming (SLSQP) was presented in [26]. It is easy to see that, depending on the number of hyperparameters per model and the number of models, this approach may be difficult to efficiently scale.
In this work, a bi-level optimization strategy is chosen to solve Equation (10) where $\underline{\beta}$ is optimized on the outer loop with the inner loop performing sample allocation optimization for each instance of $\underline{\beta}$:
@@
Here, it is emphasized that the inner loop optimization problem depends on $\underline{\beta}$ through the required estimates of model correlations $\underline{\rho}(\underline{\beta})$ and $\mathbf{P}(\underline{\beta})$. The bi-level approach decouples the model tuning problem, allowing different off-the-shelf optimization algorithms for the outer level to be combined with varying ACV estimator schemes (Table 1) on the inner level. This includes the parametric ACV approach for an additional search over model graphs on the inner level as demonstrated later on in the manuscript. Furthermore, the outer level optimization strategy only requires a single surrogate for the $\mathbb{V}ar\!\left[\tilde{Q}\left(\underline{\beta},\underline{\mathbf{z}}\right)\right]$ resulting from the inner level optimization, rather than a surrogate for each term of $\underline{\rho}(\underline{\beta})$ and $\mathbf{P}(\underline{\beta})$ as in the case of the all-at-once optimization.
@@
The primary challenge of the bi-level approach is that the outer-loop objective function is noisy since $\mathbb{V}ar\!\left[\tilde{Q}\right$] is evaluated from estimates of $\underline{\rho}(\underline{\beta})$ and $\mathbf{P}(\underline{\beta})$ based on finite (and often limited) pilot sampling. Note that either online pilot or pilot projection can be used to solve the inner-level sample allocation optimization depending on whether the accuracy or efficiency of the solution will be prioritized, respectively. Pilot projection is adopted in this work in order to avoid over-investment in the model tuning process, since only the samples for $\underline{\beta}^{*}$ can be reused.
@@
For the outer loop, the efficient global optimization (EGO) algorithm [40] is used for its ability to handle the effects of numerical noise while minimizing $\mathbb{V}ar\!\left[\tilde{Q}\right$]. EGO is a global surrogate-based optimizer that approximates the objective function with a Gaussian process (GP) response surface (i.e., a Bayesian optimization technique). EGO fits an initial GP to objective function values evaluated at $N_{init}$ random candidate points [^7] This work uses $N_{init}=(N_{\beta}+1)(N_{\beta}+2)/2$ for all examples, where $N_{\beta}=|\underline{\beta}|$ is the total number of tunable hyperparameters. and then iteratively refines the GP using points selected from the maximization of an expected improvement (EI) acquisition function. A strength of EGO is its ability to select the next candidate point in such a way that balances the need to minimize the approximate objective (exploitation) with the need to improve the approximation where it is most uncertain (exploration). Algorithm 1 summarizes the use of EGO for solving the model tuning optimization problem.Note that while EGO is an effective choice for low-dimensional model tuning problems like those presented in Section 4, a more scalable optimization algorithm (e.g., trust-region model management [41]) is recommended for locally-optimal solution to higher dimensional problems [28].
@@
In this work, the portion of the overall computational budget, $C_{budget}$, dedicated to model tuning is loosely controlled by specifying the number of pilot samples, $N_{pilot}$, and maximum number of iterations, $N_{iter}$ for the EGO algorithm. It is important to point out that a potentially significant portion of computation performed during model tuning can be reused in a subsequent multifidelity uncertainty propagation process. For example, $N_{pilot}$ evaluations of the high-fidelity model can be saved and reused as well as the evaluations of the low-fidelity models at the optimal hyperparameters, $\underline{\beta}^{*}$. Note that using a convergence tolerance that terminates EGO based on diminishing returns from model tuning could be used in lieu of (or in addition to) setting $N_{iter}$. Doing so would have some practical advantages, but was not pursued here and instead left for future studies.
@@
where it is again assumed that all $M$ low-fidelity models are tunable. Thus, when subsequently constructing the ACV estimator, the overall computational budget is modified as $C_{budget}-C_{\mathcal{T}}$.
@@
1:Input: Number of initial candidate points $N_{init}$, number of iterations $N_{iter}$, number of pilot samples $N_{pilot}$, computational budget $C_{budget}$
2:Let: $J(\underline{\beta})\equiv\min_{\underline{\mathbf{z}}}\mathbb{V}ar\!\left[\tilde{Q}(\underline{\beta},\underline{\mathbf{z}})\right]\quad\text{s.t.}\quad\mathcal{C}(\underline{\beta},\underline{\mathbf{z}})<C_{budget}$ $\triangleright$ EGO objective function
3:Randomly select $N_{init}$ hyperparameter candidate values $\{\underline{\beta}_{i}\}_{i=1}^{N_{init}}$
4:Initialize dataset: $\mathcal{D}_{0}=\{\underline{\beta}_{i},J(\underline{\beta}_{i})\}_{i=1}^{N_{init}}$ $\triangleright$ Evaluate $J(\underline{\beta})$ for each candidate
5:for $t=1$ to $N_{iter}$ do
6:  Fit Gaussian Process surrogate model to $\mathcal{D}_{t-1}$
7:  Select candidate point $\underline{\beta}_{t}$ to maximize the expected improvement in $J$
8:  Estimate $\underline{\rho}(\underline{\beta}_{t})$ and $\mathbf{P}(\underline{\beta}_{t})$ with $N_{pilot}$ pilot samples
9:  Update dataset: $\mathcal{D}_{t}=\mathcal{D}_{t-1}\cup\{\underline{\beta}_{t},J(\underline{\beta}_{t})\}$
10:end for
11:Return: Optimal hyperparameters $\underline{\beta}^{*}=\arg\min_{(\underline{\beta}_{i},J(\underline{\beta}_{i}))\in\mathcal{D}_{N_{iter}}}J(\underline{\beta}_{i})$
2.6 Summary: Online Multifidelity Estimators with Model Tuning and Iterated ACV
Figure 1 summarizes the overall procedure for obtaining ACV estimators, $\tilde{Q}\left(\underline{\beta},\underline{\alpha},\underline{\mathbf{z}}\right)$, via Equation (11) in the practical setting where neither optimal low-fidelity model settings ($\underline{\beta}$) nor knowledge of model statistics ($\mathbf{P}$, $\underline{\rho}$) are available a priori. The diagram highlights two instances where computations can be reused from the previous step to improve efficiency: 1) Iterated ACV reuses pilot samples from the model tuning step to form initial estimates of $\mathbf{P}$ and $\underline{\rho}$ that are then iterated upon, and 2) the sample allocation build out starts from the final collection of shared pilot samples from the iterated ACV step, only needing to evaluate the low-fidelity models at samples not shared across all models. See Algorithm 2 for more details.
Figure 1: The online procedure for constructing a multifidelity estimator using automated model tuning and iterated ACV used in this work.
@@ 2.6 Summary: Online Multifidelity Estimators with Model Tuning and Iterated ACV
Algorithm 2 Online Multifidelity Estimator with Model Tuning and Iterated ACV
1:Input: Computational budget $C_{budget}$, number of initial candidate hyperparameter values $N_{init}$, number of optimization iterations $N_{iter}$, number of (initial) pilot samples $N_{pilot}$, bounds for low-fidelity hyperparameters $\underline{\beta}$
2:
3:Step 1: Solve model tuning optimization problem $\triangleright$ Via Algorithm 1
4:     Input: $N_{iter}$, $N_{pilot}$, $N_{init}$, bounds for $\underline{\beta}$
5:     Return: Optimal hyperparameters $\underline{\beta}^{*}$, model tuning overhead $C_{\mathcal{T}}$
@@
7:Step 2: Perform iterated ACV sample allocation optimization
8:     Input: $\underline{\beta}^{*}$, remaining budget $C_{budget}-C_{\mathcal{T}}$
9:     Return: Optimal sample allocation $\underline{\mathbf{z}}^{*}$, estimated model statistics $\{\mathbf{P},\underline{\rho}\}$, estimated variance $\mathbb{V}ar\!\left[\tilde{Q}(\underline{\beta}^{*},\underline{\mathbf{z}}^{*})\right]$
@@
11:Step 3: Evaluate models according to optimal sample allocation
12:     Input: $\underline{\beta}^{*}$, $\underline{\mathbf{z}}^{*}$
13:     Return: Outputs for all models evaluated on $\underline{\mathbf{z}}^{*}$
@@
15:Step 4: Form ACV estimator $\triangleright$ Via Equation (11)
16:     Input: Sample allocation model outputs, model statistics $\{\mathbf{P},\underline{\rho}\}$
17:     Return: ACV estimator $\tilde{Q}$
18:
19:Return: Optimal hyperparameters $\underline{\beta}^{*}$, ACV estimator $\tilde{Q}$, estimated variance $\mathbb{V}ar\!\left[\tilde{Q}(\underline{\beta}^{*},\underline{\mathbf{z}}^{*})\right]$